# SwarnRaft: Leveraging Consensus for Robust Drone Swarm Coordination in GNSS-Degraded Environments

### v1

@@ III-E1 Local Vote Broadcast
Based on the threshold test, peer $j$ issues a binary vote
$v^{(j)}_{i,k}=\begin{cases}+1,&e^{(j)}_{i,k}\leq T,\\[6.0pt] -1,&e^{(j)}_{i,k}>T.\end{cases}$
A vote of +1+1+ 1 indicates honest, while 1-1- 1 signals faulty. This local decision is then broadcast to all other peers, forming the input for the subsequent SwarmRaft stage.
Leader computes votes for each node and flags their GNSS as honest or faulty, i.e., compute $S_{i}=\sum_{j}v^{(j)}_{i,k}$:
  • If $S_{i}\geq n-f$, decide $i$ is honest.
  • If $S_{i}<-(n-f)$, decide $i$ is faulty.
@@ III-F Position Correction & Recovery
If decision $i$ is is faulty, the position is set to
$\tilde{\mathbf{x}}_{i,k}=\mathrm{median}_{j}\bigl{\{}\hat{\mathbf{x}}^{(j)}_{i,k}\bigr{\}}.$
Node $i$ resets its INS state to $\tilde{\mathbf{x}}_{i,k}$; otherwise it retains its fused estimate.
Note: With a low probability we can flag one or several honest sensors as faulty, which may lead to a majority of the sensors flagged as faulty for the current $k$. In such a case, we do not compute the median during the recovery phase, but recover based on INS only.
@@ III-G Security & Fault-Tolerance Analysis
Integrity: All inter-node communications are authenticated using pre-distributed cryptographic keys, preventing forgery and tampering. Consensus-based validation ensures the trustworthiness of the the reported data. Although rare statistical anomalies can incorrectly flag an honest sensor as faulty, such false positives are mitigated by majority voting. In edge cases where too many nodes are flagged, the system falls back to inertial (INS) data for continuity, preserving mission integrity at the cost of reduced accuracy.
Communication Efficiency: SwarmRaft is designed to scale efficiently. The leader incurs a communication and computation cost of $O(n)$ per round, while non-leader nodes communicate only with the leader at $O(1)$ cost. This model minimizes bandwidth usage and processing load as the swarm scales. Raft log replication further ensures consistency across all nodes.
Fault Tolerance: The protocol tolerates up to $f$ Byzantine faults, including GNSS spoofing, range manipulation and collusion. By requiring $n\geq 2f+1$, SwarmRaft ensures that honest node votes dominate and that consensus decisions remain trustworthy even under adversarial conditions. The coordination of the leader of the Raft and the replicated logs reinforce fault-tolerant behavior during ongoing operations.
@@
Figure 1: Visualization of SwarmRaft recovery under different swarm sizes ($N=5,10,15$). Yellow stars represent recovered positions using consensus and range-based fusion.
@@ IV Prototype Implementation
To evaluate SwarmRaft under adversarial localization scenarios, we developed a modular simulation framework in Python that models both the dynamics and vulnerabilities of drone swarms. The framework simulates each drone’s ground-truth motion, inertial navigation estimates, GNSS observations (including spoofed signals), and noisy inter-drone ranging data. A dedicated leader node module performs fused position estimation, fault detection through residual analysis and majority voting, and robust recovery based on peer consensus. To support performance benchmarking, the simulator includes tools for running large-scale Monte Carlo experiments that vary swarm size and attacker configuration, producing statistical metrics including MAE. Additionally, the framework supports real-time visualization and post-simulation plotting to analyze error evolution and swarm behavior.
Each drone instance maintains a ground-truth position vector $\mathbf{x}_{i}^{\text{true}}$, a GNSS-based measurement $\mathbf{z}_{i}$, and an inertial estimate $\mathbf{x}_{i}^{\text{INS}}$ defined by:
$\mathbf{x}_{i}^{\text{INS}}=\mathbf{x}_{i}^{\text{prev}}+\Delta\mathbf{u}_{i}+\boldsymbol{\varepsilon}_{\text{INS}},\quad\mathbf{z}_{i}=\mathbf{x}_{i}^{\text{true}}+\boldsymbol{\varepsilon}_{\text{GNSS}}.$ (1)
Inter-agent fusion is performed via weighted averaging of local and remote estimates. The fused position from agent $j$ to agent $i$ is given by:
$\hat{\mathbf{x}}_{i}^{(j)}=\alpha\mathbf{z}_{i}+(1-\alpha)\left(\mathbf{x}_{j}^{\text{INS}}+\frac{d_{ij}\cdot(\mathbf{x}_{i}^{\text{prev}}-\mathbf{x}_{j}^{\text{INS}})}{\|\mathbf{x}_{i}^{\text{prev}}-\mathbf{x}_{j}^{\text{INS}}\|}\right).$ (2)
The weighting coefficient $\alpha$ depends on the variance of GNSS and range measurements:
$\alpha=\frac{\sigma_{r}^{2}}{\sigma_{r}^{2}+\sigma_{z}^{2}}.$ (3)
Each agent compares the fused estimate $\hat{\mathbf{x}}_{i}^{(j)}$ to its own GNSS reading. If the residual is below a threshold $\tau$, a positive vote is cast; otherwise, the vote is negative:
$v_{ij}=\begin{cases}+1,&\text{if }\|\hat{\mathbf{x}}_{i}^{(j)}-\mathbf{z}_{i}\|\leq\tau\\ -1,&\text{otherwise}\end{cases}.$ (4)
Faulty agents are identified when the number of negative votes exceeds the trust threshold $(n-f)$, where $n$ is the swarm size and $f$ is the maximum number of faults considered. Position recovery is performed by aggregating fused estimates from peers using the coordinate-wise median:
$\hat{\mathbf{x}}_{i}=\operatorname{median}\left\{\hat{\mathbf{x}}_{i}^{(j)}\;\middle|\;j\in\mathcal{N}(i)\right\}.$ (5)
@@
The prototype accommodates configurable attack injection, including GNSS spoofing, range distortion, and randomized attacker behavior. The implementation serves as the experimental foundation for the visual and quantitative evaluations discussed in subsequent sections. The source code for SwarnRaft is publicly available on GitHub [39].
V Numerical experiments
@@ V-A Qualitative Results for Position Recovery
Figure 1 illustrates the SwarmRaft consensus-based recovery mechanism across varying swarm sizes $N\in\{5,10,15\}$, simulating aerial drone localization using GNSS and INS. Each subfigure demonstrates how noisy or spoofed measurements are corrected through distributed voting and recovery. This visualization not only showcases concrete recovery behavior but also reveals the spatial structure of localization errors in both small- and large-scale swarms, complementing the accompanying statistical evaluation.
@@
  • Green circle ( $\circ$): the ground-truth position $\mathbf{x}_{i}^{\mathrm{true}}$.
  • Blue cross ( $\times$): the GNSS measurement $\mathbf{z}_{i}=\mathbf{x}_{i}^{\mathrm{true}}+\boldsymbol{\varepsilon}_{i}^{\mathrm{GNSS}}$, where $\boldsymbol{\varepsilon}_{i}^{\mathrm{GNSS}}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$.
  • Red triangle ( $\triangle$): the INS estimate $\mathbf{x}_{i}^{\mathrm{INS}}$, which accumulates integration errors and noise.
  • Yellow star ( $\star$): the recovered position $\hat{\mathbf{x}}_{i}$ computed via the SwarmRaft algorithm.
The SwarmRaft protocol fuses inter-drone distance measurements and local INS readings using a consensus scheme. Each node receives fused position suggestions from its neighbors and performs majority voting to detect inconsistencies. If a node is suspected to be compromised, its position is recovered via a robust aggregation (e.g., coordinate-wise median) of neighbor estimates.
Mathematically, the fused estimate from a neighbor $j$ to a node $i$ is computed as:
$\hat{\mathbf{x}}_{i}^{(j)}=\alpha\mathbf{z}_{i}+(1-\alpha)\left(\mathbf{x}_{j}^{\mathrm{INS}}+\frac{d_{ji}\cdot(\mathbf{x}_{i}^{\mathrm{prev}}-\mathbf{x}_{j}^{\mathrm{INS}})}{\|\mathbf{x}_{i}^{\mathrm{prev}}-\mathbf{x}_{j}^{\mathrm{INS}}\|}\right).$ (6)
Here, $\alpha=\frac{\sigma_{r}^{2}}{\sigma_{r}^{2}+\sigma_{\mathrm{GNSS}}^{2}}$ is a weighting coefficient derived from error variances, and $d_{ji}$ is the noisy range measurement between drones $j$ and $i$.
@@ V-B Quantitative Evaluation Across Swarm Sizes and Attacks
To complement the visual analysis of SwarmRaft’s behavior, we conduct a large-scale statistical evaluation using simulated drone swarms with varying sizes and attack conditions. For each swarm size $N\in\{5,10,15\}$ and each number of attacked drones $f\in\{1,2,\dots,N-1\}$, we run 10 Monte Carlo trials. During each trial, $f$ drones are randomly selected to be compromised via GNSS spoofing and range manipulation. The remaining drones follow the nominal sensing model, where the GNSS and INS errors are drawn from zero-mean Gaussian distributions.
Each Monte Carlo trial produces two sets of localization error metrics:
  • GNSS Baseline Error: The mean absolute error (MAE) is calculated based on the raw GNSS readings:
    $\mathrm{MAE}_{\mathrm{GNSS}}=\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{z}_{i}-\mathbf{x}_{i}^{\mathrm{true}}\|.$
  • SwarmRaft-Recovered Error: The same metrics are computed using the recovered positions $\hat{\mathbf{x}}_{i}$ after consensus-based correction:
    $\mathrm{MAE}_{\mathrm{SwarmRaft}}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{\mathbf{x}}_{i}-\mathbf{x}_{i}^{\mathrm{true}}\|.$
The results are aggregated across all trials and shown in Figure 2. Each curve depicts the average error as a function of the number of attacked drones $f$, grouped by the total swarm size $N$.
Figure 2: Error trends for GNSS and SwarmRaft-recovered positions under varying attack intensities and swarm sizes. Each curve is averaged over 10 trials.
@@
    GNSS-only estimates degrade significantly as $f$ increases, particularly in larger swarms.
@@
    The SwarmRaft algorithm consistently reduces MAE compared to raw GNSS readings, even when a majority of drones are attacked.
@@
    The recovered errors remain stable as swarm size increases, suggesting scalability and resilience of the consensus mechanism.
These results demonstrate that SwarmRaft is capable of robust fault detection and position recovery even under adversarial and large-scale conditions.
Figure 3: Boxplot of mean absolute error (MAE) for GNSS-only positioning and SwarmRaft-recovered estimates across all simulation trials. The recovered positions exhibit lower median error and reduced variance, indicating improved accuracy and robustness under adversarial conditions.
(a) $N=5$
(b) $N=10$
(c) $N=15$
Figure 4: Comparison of mean and median absolute localization errors for GNSS and SwarmRaft-recovered estimates under varying numbers of attacked drones. Each subfigure corresponds to a different swarm size. Dashed lines represent means; solid lines denote medians. SwarmRaft reduces sensitivity to outliers, especially in larger swarms.
Figure 3 compares the distribution of localization errors under GNSS-only and SwarmRaft-recovered positioning across diverse simulation trials. The GNSS error distribution shows a higher median, wider spread, and more extreme values, reflecting its vulnerability to spoofing and sensor faults. In contrast, the SwarmRaft-corrected estimates demonstrate a substantially reduced median error and tighter interquartile range, with fewer outliers. This indicates both improved accuracy and robustness, underscoring the protocol’s effectiveness in mitigating adversarial localization noise through consensus and peer fusion.
V-C Comparison of Mean and Median Errors
Figures 4 present the evolution of absolute localization errors as the number of compromised agents increases, across swarm sizes $N=5$, 101010, and 151515. Each subplot compares GNSS-only positioning with estimates recovered using the SwarmRaft protocol.
GNSS results exhibit a growing discrepancy between mean and median errors as more nodes are spoofed, revealing vulnerability to extreme outliers. In contrast, the SwarmRaft-based estimates maintain a smaller gap between these statistics, indicating robustness to adversarial noise. This behavior suggests that the consensus and fusion mechanism is capable of filtering out anomalous contributions even in highly degraded conditions. Additionally, as the swarm size increases, the recovered medians become increasingly stable, reflecting the benefits of larger peer sets in collaborative state estimation.
VI Discussion
The experiments in Section V evaluate the practical behavior and performance of the proposed SwarmRaft protocol under controlled adversarial conditions, including GNSS spoofing and range measurement perturbations. Simulations were performed in varying swarm sizes and numbers of compromised nodes.
Figure 1 visually demonstrates SwarmRaft’s ability to detect and isolate faulty agents, even in small swarms with limited redundancy. The recovered positions align closely with the truth of the ground, highlighting the effectiveness of peer triangulation and consistency-based voting. The quantitative results in Figures 3 and 4 demonstrate that SwarmRaft significantly reduces both mean and median localization errors, with accuracy improving as the size of the swarm increases. As more peers become available, the system maintains low median errors even in the presence of multiple compromised nodes. The corresponding boxplot further confirms the robustness of the method, showing reduced variance and fewer outliers under adversarial conditions.
These findings highlight the core strengths of SwarmRaft, including its robustness to localized spoofing without requiring prior attacker identification, its scalability in decentralized deployments with varying trust levels, and its adaptability to sensor noise and signal degradation. The protocol operates without a centralized fusion authority and tolerates Byzantine behavior under mild density assumptions, making it well-suited for distributed robotic swarms operating in partially adversarial environments.
Nonetheless, several limitations remain. The current design assumes bounded noise and reliable relative ranging, which may degrade in GNSS-denied urban canyons or under active jamming. The protocol also treats all peers equally, without incorporating trust weighting or historical behavior. Moreover, while efficient in simulation, real-time deployment may face latency and packet loss challenges that affect performance.
VII Conclusion
This paper presented SwarmRaft, a consensus-based protocol for resilient UAV swarm localization in GNSS-degraded and adversarial environments. By fusing inertial measurements with peer-to-peer ranging and validating positional consistency through distributed voting, SwarmRaft achieves accurate state estimation even in the presence of GNSS spoofing and range manipulation. In contrast to centralized or purely sensor-driven approaches, it integrates fault detection, consensus-based validation, and position recovery into a lightweight and scalable framework inspired by Raft.
Extensive simulations across varying swarm sizes and attack conditions demonstrated that SwarmRaft consistently reduces both mean and median localization errors relative to raw GNSS measurements. Notably, the protocol maintains stable performance even as the number of compromised agents increases, with larger swarms enhancing robustness through increased redundancy and more reliable peer estimates.
Beyond adversarial fault mitigation, SwarmRaft offers a generalizable architecture for collaborative localization under uncertainty, requiring minimal assumptions about trust, infrastructure, or communication hierarchy. Its compatibility with standard UAV sensing modalities (GNSS, INS, and ranging) and its ability to operate in fully decentralized settings support its applicability in real-world missions such as disaster response, infrastructure inspection, and autonomous logistics.
Future work will focus on extending the protocol with dynamic trust modeling, adaptive voting thresholds, and confidence-weighted fusion to enhance resilience against sensor drift and coordinated adversarial behavior. In addition, performance under asynchronous communication and packet loss should be evaluated, alongside validation on physical UAV swarms operating in real-world conditions. Ultimately, SwarmRaft establishes a principled foundation for robust, autonomous multi-agent systems capable of reliable operation in contested and degraded environments.

### v2

@@ III-E1 Local Vote Broadcast
Based on the threshold test, the leader assigns a binary decision for each node $i$:
$v_{i,k}=\begin{cases}+1,&e_{i,k}\leq T,\\[6.0pt] -1,&e_{i,k}>T.\end{cases}$
A decision of +1+1 indicates that node $i$’s reported position is consistent with the probabilistic regions derived from neighbor constraints, while 1-1 signals a potential fault. These decisions are then broadcast to all nodes, forming the input for the subsequent SwarmRaft stage.
  • If $S_{i}\geq 0$, decide $i$ is honest.
  • If $S_{i}<0$, decide $i$ is faulty.
@@ III-F Position Correction & Recovery
Once a node $i$ is flagged as faulty, its position is recomputed using only the non-faulty neighbors’ information. Let $\mathcal{N}_{i}^{\text{good}}$ denote the set of neighbors of node $i$ that are not flagged as faulty. Each neighbor $j\in\mathcal{N}_{i}^{\text{good}}$ provides a reported position $\mathbf{x}_{j,k}$ and the measured distance $d_{i,j}$ to node $i$.
We model the feasible location of the faulty node $i$ as the intersection of spheres centered at each non-faulty neighbor with radius equal to the measured distance:
$\mathcal{S}_{j}=\bigl\{\mathbf{x}\in\mathbb{R}^{3}\;|\;\|\mathbf{x}-\mathbf{x}_{j,k}\|=d_{i,j}\bigr\}.$
The intersection of these spheres defines the feasible region for the node’s true position.
To compute a single position estimate $\tilde{\mathbf{x}}_{i,k}$, we solve a nonlinear least-squares problem that minimizes the squared deviations from all spheres:
$\tilde{\mathbf{x}}_{i,k}=\arg\min_{\mathbf{x}\in\mathbb{R}^{3}}\sum_{j\in\mathcal{N}_{i}^{\text{good}}}\bigl(\|\mathbf{x}-\mathbf{x}_{j,k}\|-d_{i,j}\bigr)^{2}.$
This formulation finds the point that best fits all range measurements from non-faulty neighbors, effectively performing a multilateration.
In practice, we initialize the optimization at the centroid of the non-faulty neighbors:
$\mathbf{x}^{(0)}_{i,k}=\frac{1}{|\mathcal{N}_{i}^{\text{good}}|}\sum_{j\in\mathcal{N}_{i}^{\text{good}}}\mathbf{x}_{j,k},$
and refine $\tilde{\mathbf{x}}_{i,k}$ via iterative least-squares (e.g., using a soft-$L_{1}$ loss to reduce sensitivity to residual errors). Once obtained, the node resets its INS state to $\tilde{\mathbf{x}}_{i,k}$.
Note: If too few non-faulty neighbors remain (e.g., due to false positives), we skip multilateration and rely solely on the node’s INS propagation. This ensures that the recovery procedure remains robust even under adversarial conditions.
Note: In rare cases, multiple honest sensors may be falsely flagged as faulty, potentially leaving too few verified neighbors for multilateration. When this occurs, we skip median/multilateration recovery and rely solely on the node’s INS propagation to update its state.
@@ III-G Security & Fault-Tolerance Analysis
Integrity: All inter-node communications are authenticated using pre-distributed cryptographic keys, preventing forgery and tampering. Consensus-based validation ensures the trustworthiness of the reported data. Although rare statistical anomalies can incorrectly flag an honest sensor as faulty, such false positives are mitigated by majority voting. In edge cases where too many nodes are flagged, the system falls back to inertial (INS) data for continuity, preserving mission integrity at the cost of reduced accuracy.
Communication Efficiency: SwarmRaft is designed to scale efficiently. The leader incurs a communication and computation cost of O(2(n1))O(2(n-1)) per round, while non-leader nodes communicate only with the leader at $O(2)$ cost. This model minimizes bandwidth usage and processing load as the swarm scales. Raft log replication further ensures consistency across all nodes.
Fault Tolerance: The protocol tolerates up to $f$ Byzantine faults, including GNSS spoofing, range manipulation, and collusion. By requiring $n\geq 2f+1$, SwarmRaft ensures that honest node votes dominate and that consensus decisions remain trustworthy even under adversarial conditions. The coordination of the leader of the Raft and the replicated logs reinforces fault-tolerant behavior during ongoing operations.
@@
Figure 1: Visualization of SwarmRaft recovery under different swarm sizes ($N=6,11,16$). Yellow stars represent recovered positions using consensus and range-based fusion.
@@ IV Prototype Implementation
To evaluate SwarmRaft under adversarial localization scenarios, we developed a modular simulation framework in Python that models both the dynamics and vulnerabilities of drone swarms. The framework simulates each drone’s ground-truth motion, inertial navigation estimates, GNSS observations (including spoofed signals), and noisy inter-drone ranging data. A dedicated leader node module performs position verification, fault detection through voting, and recovery via iterative multilateration. To support performance benchmarking, the simulator includes tools for running large-scale Monte Carlo experiments that vary swarm size and attacker configuration, producing statistical metrics including mean absolute error (MAE). Additionally, the framework supports real-time visualization and post-simulation plotting to analyze error evolution and swarm behavior.
IV-A Algorithmic Procedure
The leader-based verification algorithm operates in two main stages. First, each drone’s reported position is cross-validated against inter-drone distance measurements. The leader tallies votes from all pairwise checks, where consistent neighbors contribute positive votes and inconsistent ones contribute negative votes. Any drone that accumulates a majority of negative votes is flagged as potentially faulty.
In the second stage, the leader refines the positions of flagged drones using multilateration. This involves solving a least-squares problem with respect to verified neighbors, iteratively adjusting the estimated position until convergence or until a maximum number of iterations is reached. If the corrected estimate deviates significantly from the originally reported position, the corrected value replaces the spoofed report; otherwise, the drone is reclassified as valid. This process guarantees that up to $f$ malicious or faulty drones can be corrected in a swarm of size $n\geq 2f+1$.
Algorithm 1 Leader-based Neighbor Verification with Voting and Multilateration
1:Reported positions $\mathbf{x}^{\rm rep}_{i}$ from all UAVs $i=1,\dots,n$ Inter-UAV distance matrix $D=[d_{ij}]$ measured from true geometry Base tolerance $\epsilon>0$, maximum iterations $k_{\max}$
2:Verified positions $\mathbf{x}^{\rm ver}_{i}$ and fault flags for all UAVs
3:Leader collects $\mathbf{x}^{\rm rep}_{1:n}$ and distance matrix $D$
4:for each UAV $A=1$ to $n$ do $\triangleright$ Stage 1: Voting
5:  Initialize vote counter $v_{A}\leftarrow 0$
6:  for each neighbor $B\neq A$ do
7:   Compute distance from reports: $\hat{d}_{AB}=\|\mathbf{x}^{\rm rep}_{A}-\mathbf{x}^{\rm rep}_{B}\|$
8:   if $|\hat{d}_{AB}-d_{AB}|<\tau$ then $\triangleright$ consistency check
9:     $v_{A}\leftarrow v_{A}+1$
10:   else
11:     $v_{A}\leftarrow v_{A}-1$
12:   end if
13:  end for
14:  Flag AA as faulty if $v_{A}<0$
15:  Initialize $\mathbf{x}^{\rm ver}_{A}\leftarrow\mathbf{x}^{\rm rep}_{A}$
16:end for
17:for each UAV AA flagged as faulty do $\triangleright$ Stage 2: Multilateration refinement
18:  Select non-faulty neighbors $\mathcal{N}_{A}$
19:  if $|\mathcal{N}_{A}|<3$ then
20:   $\mathcal{N}_{A}\leftarrow$ all other UAVs $\triangleright$ fallback if insufficient anchors
21:  end if
22:  Initialize estimate $\mathbf{p}\leftarrow\mathbf{x}^{\rm rep}_{A}$
23:  for $k=1$ to $k_{\max}$ do
24:   Update $\mathbf{p}$ by solving least-squares residuals:
$\mathbf{p}\leftarrow\arg\min_{\mathbf{q}\in\mathbb{R}^{3}}\sum_{j\in\mathcal{N}_{A}}\big(\|\mathbf{q}-\mathbf{x}^{\rm rep}_{j}\|-d_{jA}\big)^{2}$
25:   if converged then
26:     break
27:   end if
28:  end for
29:  Compute deviation $\Delta_{A}=\|\mathbf{x}^{\rm rep}_{A}-\mathbf{p}\|$
30:  if $\Delta_{A}>\epsilon$ then
31:   $\mathbf{x}^{\rm ver}_{A}\leftarrow\mathbf{p}$ $\triangleright$ replace spoofed report
32:  else
33:   $\mathbf{x}^{\rm ver}_{A}\leftarrow\mathbf{x}^{\rm rep}_{A}$
34:   Mark AA as non-faulty
35:  end if
36:end for
37:return Verified positions $\mathbf{x}^{\rm ver}_{1:n}$ and fault flags
@@ IV-A Algorithmic Procedure
The prototype accommodates configurable attack injection, including GNSS spoofing, range distortion, and randomized attacker behavior. The implementation serves as the experimental foundation for the visual and quantitative evaluations discussed in subsequent sections. The source code for SwarmRaft is publicly available on GitHub [43].
V Numerical experiments and results
@@ V-A Qualitative Results for Position Recovery
Figure 1 illustrates the SwarmRaft recovery mechanism across varying swarm sizes. The three subplots correspond to $(n,f)\in\{(6,2),(11,5),(16,7)\}$, i.e., swarms with two, five, and seven faulty drones, respectively. Each visualization demonstrates that SwarmRaft is able to correctly identify spoofed drones and restore their reported positions to match ground truth via multilateration. For spoofed drones, the recovered position (yellow star) aligns with the true position (green cross), while the spoofed report (red triangle) is visibly displaced. Arrows further highlight the deviation between true and spoofed reports, making the correction visually evident. SwarmRaft’s ability to detect and isolate faulty agents, even in small swarms with limited redundancy. The recovered positions align closely with the truth of the ground, highlighting the effectiveness of peer triangulation and consistency-based voting. Each subfigure demonstrates how noisy or spoofed measurements are corrected through distributed voting and recovery. This visualization not only showcases concrete recovery behavior but also reveals the spatial structure of localization errors in both small- and large-scale swarms, complementing the accompanying statistical evaluation.
@@
  • Red triangle ($\blacktriangle$): the spoofed reported position $\mathbf{x}_{i}^{\mathrm{rep}}$.
  • Yellow star ($\star$): the recovered position $\hat{\mathbf{x}}_{i}$ computed via multilateration.
The SwarmRaft protocol verifies positions through two stages. First, consistency voting checks whether reported inter-drone distances match measured distances; drones that fail majority voting are flagged as faulty. Second, flagged drones undergo multilateration using non-faulty neighbors as anchors. If the corrected estimate significantly deviates from the reported position, the spoofed report is replaced with the multilateration-based recovery. This ensures that up to $f$ compromised drones can be tolerated in a swarm of size $n\geq 2f+1$.
@@ V-B Quantitative Evaluation Across Swarm Sizes and Attacks
To complement the qualitative visualizations, we conduct a large-scale statistical evaluation of SwarmRaft under varying swarm sizes and adversarial conditions. For each swarm size $n\in\{3,5,7,9,11,13,15,17\}$ and each number of attacked drones $f\in\{1,2,\dots,8\}$, we run Monte Carlo trials where $f$ drones are randomly selected as spoofed. Spoofed drones report displaced positions and manipulated ranges, while the remaining drones report correctly.
Each trial produces two error metrics:
  • Baseline Error: The MAE of reported (potentially spoofed) positions relative to ground truth.
  • Recovered Error: The MAE of the recovered and multilaterated positions $\hat{\mathbf{x}}_{i}$ produced by SwarmRaft.
Figure 2 shows the scaling performance of SwarmRaft as swarm size grows. For a given number of faulty drones $f$, we consider the minimal swarm size $n=2f+1$, which guarantees that the majority of drones remain non-faulty. The plot reports the mean recovery error on a logarithmic scale, averaged over $10,000$ Monte Carlo simulations. As the size of the swarm increases relative to $f$, the recovery accuracy improves dramatically, with a mean error dropping from approximately 19 m to 0.28 m.
Figure 2: Scaling performance of SwarmRaft as swarm size increases. Mean recovery error (log scale) decreases significantly with larger swarms, from  19 m to  0.28 m.
@@
    Reported (spoofed) positions quickly diverge from ground truth, reflecting the impact of adversarial manipulation.
@@
    SwarmRaft consistently reduces error compared to the baseline, even when nearly half of the swarm is compromised.
@@
    Recovery accuracy improves with swarm size, demonstrating scalability and resilience of the protocol.
System Relevant dimension Complexity per update Typical scale
GNSS (LS solution) $m\approx 6$1010 satellites $\mathcal{O}(m^{3})$ (constant) Few $10^{2}$ ops
INS propagation $n\approx 9$1515 states $\mathcal{O}(n)$ Few $10^{2}$ ops
GNSS+INS (Kalman) $n\approx 15$ states $\mathcal{O}(n^{3})$ $\sim$3,000 ops
Proposed algorithm NN drones Regular: $\mathcal{O}(N)$, Leader: $\mathcal{O}(N^{2})$ 2002006,0006{,}000 ops (for $N\leq 17$)
TABLE II: Computational complexity comparison of GNSS, INS, and the proposed verification algorithm.
Figure 2 further compares error distributions across trials. Baseline errors show high spread, large medians, and many outliers, confirming vulnerability to spoofing. In contrast, SwarmRaft-corrected positions exhibit low medians, tight interquartile ranges, and few outliers, reflecting robustness under adversarial conditions.
V-C Computational Complexity
To provide a fair comparison, we summarize the computational complexity of the proposed verification algorithm against conventional GNSS and INS solutions. For GNSS, position computation after signal tracking reduces to solving a set of pseudorange equations, typically involving $m\approx 6$1010 satellites. This least-squares solution requires a matrix inversion of size $m\times m$, with complexity $\mathcal{O}(m^{3})$, which is effectively constant in practice. INS propagation involves simple state integration (position, velocity, attitude), with complexity $\mathcal{O}(n)$ for a state dimension $n\approx 9$1515. When GNSS and INS are fused using a Kalman filter, the update step has complexity $\mathcal{O}(n^{3})$ due to covariance updates, where $n\approx 15$ states, yielding on the order of 3,0003{,}000 arithmetic operations per update.
In contrast, in our proposed distributed swarm verification scheme, each regular node computes distances to all other nodes, yielding $\mathcal{O}(N)$ operations per update for swarm size $N$. The leader node, in the worst-case solvable scenario ($f\approx(N-1)/2$ faulty nodes), must iteratively correct $f$ faulty positions, resulting in $\mathcal{O}(fN)\approx\mathcal{O}(N^{2})$. For typical values ($N\leq 17$), this corresponds to fewer than 6,0006{,}000 floating-point operations, which is of the same order as a GNSS+INS Kalman update and easily handled by modern embedded processors. A summary of these computational complexities, together with the proposed SwarmRaft verification algorithm, is provided in Table II for ease of comparison.
VI Conclusion
This paper introduced SwarmRaft, a novel consensus-based protocol for resilient and decentralized UAV swarm localization in GNSS-degraded and adversarial environments. Unlike prior approaches that rely on centralized control or computationally intensive BFT protocols, SwarmRaft adapted the Raft consensus algorithm to support lightweight, crash-tolerant decision-making within resource-constrained aerial swarms. The system fused inertial navigation data with inter-drone ranging and applied distributed voting to detect and recover from sensor faults or spoofed position reports. Through extensive simulations under varying swarm sizes and attack intensities, SwarmRaft demonstrated significant improvements in both mean and median localization accuracy compared to GNSS-only baselines, particularly in scenarios with partial compromise. The protocol’s ability to maintain collective situational awareness without a central authority highlights its practical deployability for time-sensitive missions such as autonomous logistics, disaster response, and infrastructure inspection. Looking ahead, future research will focus on incorporating confidence-weighted fusion, dynamic trust scores, and asynchronous consensus strategies, as well as validating the system on physical UAV swarms in real-world settings. These advancements aim to establish SwarmRaft as a foundational architecture for fault-resilient coordination in next-generation autonomous multi-agent systems.